# KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation

### v1

@@ 4.1 Baselines
We evaluate 13 proprietary and open-source LLMs on KG-QAGen-D benchmark.
@@
GPT-4o & GPT-4o-Mini (Achiam et al., 2023; Hurst et al., 2024), Gemini-1.5-Pro & Gemini-1.5-Flash (Team et al., 2023, 2024a), and Gemini-2.0-Flash & Gemini-2.0-Flash-Lite (Team et al., 2024a).
Open-Source LLMs:
DeepSeek-V3 (Liu et al., 2024a), Llama-4-Maverick-17B-128E-Instruct-FP8 (AI, 2024a), Llama-4-Scout-17B-16E-Instruct (AI, 2024b), Llama-3.1-405B-Instruct-Turbo & Llama-3.1-70B-Instruct & Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2-72B-Instruct (Yang et al., 2024).
Given that our documents average 80k tokens, we restrict our evaluation to models supporting a context length of at least 128k tokens. Nonetheless, even long-context models require a carefully designed prompting strategy to extract relevant answers effectively (see Section D.1). We made several attempts to benchmark the recently introduced Gemini-2.5-Pro and Gemini-2.5-Flash (Google, 2025) (available in preview mode), but they consistently returned empty responses regardless of configuration. We also attempted to evaluate Gemma-2-8k (Team et al., 2024b), Mixtral-Small-32k, and Mixtral-8×22-65k (Jiang et al., 2024a), but these models frequently returned “Not found", making systematic evaluation infeasible.
For proprietary models and those accessed via Together.AI [^4] https://www.together.ai/, we utilized the LangChain [^5] https://www.langchain.com/ framework, while open-source models were employed using vLLM (Kwon et al., 2023) for efficient inference. To ensure reproducibility, we set the temperature to 0.0.
@@ 4.2 Experiment Setup
Our evaluation setup is designed to handle long financial documents and large batches of questions in a way that fits the context length constraints of modern language models. We evaluate models by prompting them with a document (or a chunk, if needed) and a batch of 50 associated questions, then measuring the quality of their answers.
As seen in Figure 3, if a document exceeds the 128K-token context limit, we split it into manageable chunks. For each chunk or full document (if it fits), we prompt the model with a batch of 50 questions associated with that segment. This design ensures consistent and scalable evaluation across documents of varying lengths. See Appendix D.1 for further details.
We use four evaluation metrics: F1 score, Edit Distance, Cosine Similarity, and LLM-as-a-Judge. Among these, F1 and LLM-as-a-Judge are highlighted in Table 4 for their complementary strengths – F1 quantifies exact token-level overlap, while LLM-as-a-Judge captures semantic correctness with human-aligned scoring. The details are given in Appendix D.2. F1 and LLM-as-a-Judge also show the highest correlation with human evaluation metrics (see Appendix D.4).
Figure 3: This figure illustrates the multi-stage process used to evaluate long-context LLMs: documents are split into chunks (if exceeding 128K tokens), paired with question batches ($\leqslant$50), and sequentially fed to the model. Model responses are collected per chunk and per question batch to ensure evaluation scalability across large, real-world documents.
@@ 4.3 Main Results
The results in Table 4 highlight a significant decline in models performance across all metrics as question complexity increases. Llama-3.1-Instruct-Turbo (405B) stands out in the Hard category, achieving an F1 score of 0.3709, significantly higher than all other models, despite not leading in Easy or Medium tasks, which could indicate stronger generalization to difficult tasks. However, it was the only model that failed to follow the requested JSON output format, requiring us to develop a separate parser. This suggests that as models become larger and more general, their ability to follow specific instructions diminishes.
Gemini-2.0-Flash (1000K) and Gemini-1.5-Flash (1000K) dominate in Easy and Medium queries but struggle on Hard, indicating a limitation in handling complex retrieval. Among open-source models, DeepSeek-V3 (671B) performs consistently but lacks dominance, while Llama-3.1-Instruct (70B) competes well on Medium, surpassing some larger proprietary models. Qwen2-Instruct (72B) underperforms across all categories, suggesting size alone does not determine effectiveness.
Table 4: This table presents the performance of 13 LLMs, evaluated across Easy, Medium, and Hard question categories. Metrics include F1 Score and LLM-as-a-Judge rating, capturing both token-level accuracy and semantic correctness. The results reveal a consistent decline in performance as question complexity increases, with notable model-specific strengths and weaknesses. denotes the models evaluated on a smaller subset due to cost constraints.
@@
Open-source LLMs
@@ 5.1 Overall Error Trends
Figure 4: Trends in error types across complexity levels. The template level is given by $P+H+\#SO$.
As question complexity increases, the LLM’s ability to retrieve and generate correct responses degrades markedly. With increase in number of hops or set operations, the percentage distribution of each error type increases (see Figure 4 for more details).
@@ 5.2 Error Analysis
Misinterpretation of Semantics LLMs frequently misinterpret task-specific terms such as “location type" or “position". Rather than grounding responses in document context, models often default to general world knowledge (e.g., answering “city" for a location type) or paraphrase the question, indicating weak comprehension of domain-specific semantics.
Implicit Information Gaps The model struggles to extract answers that are not explicitly stated but implied through document structure, such as signature blocks or formatting cues. For instance, it fails to associate a person with an organization unless the relationship is stated verbatim, ignoring visual or positional indicators that a human reader would readily infer.
Set Operation Failures To examine the impact of the number of set operations on model performance, we conducted an ablation experiment by varying one dimension within our three-dimensional parameter space while holding the others constant. Specifically, we increased the number of set operations while keeping the plurality and hop dimensions fixed. We observe that an increase in the number of set operations consistently resulted in a decline in performance metrics. Questions involving comparisons, intersections, or exclusions across entities consistently yield low F1 scores. The model fails to isolate shared or distinct roles when multiple organizations or individuals are involved, indicating limited capacity for multi-hop or set-based reasoning.
Long-Context Retrieval Errors Despite the presence of answers in the document, the model often returns “Not found", especially when relevant information appears in the middle of the context. This aligns with the “lost-in-the-middle" phenomenon (Liu et al., 2024b), where models disproportionately focus on the beginning or end of long texts, neglecting intermediate content.
@@ 6 Discussion & Conclusion
Our framework integrates structured annotations, knowledge graph construction, and templated QA extraction (Sections 3.2-3.5) to produce a dataset with finely controlled question complexity over real-world long document. This allows systematic evaluation across various dimensions with identifying nuanced model failures. Despite high scores on simple factual questions, even the strongest LLMs struggle with compositional logical inference, especially set-based comparisons and multi-hop paths, frequently failing to interpret implicit relationships and falling back on world knowledge instead of grounding answers in context (Section 5).
@@
As discussed in Sections 1 and 3.1, the ability to process long documents is essential for real-world QA tasks. Our benchmark illustrates this challenge using credit agreements, where information is densely structured, scattered, and cross-referenced. However, legal contracts, policy documents, patent filings, and other non-financial documents all exhibit similarly complex entity structures. Even in non-legal domains, such as medical records or scientific articles, multi-hop relationships and compositional multi-hop inference over entities are common, making our methodology broadly applicable.
Opportunities and Extensions
While our benchmark focuses on pre-trained LLMs, the structured design of KG-QAGen enables future adaptation, such as domain-specific fine-tuning. Expanding the dataset beyond U.S.-based financial documents could support geographically diverse evaluation.
@@ Appendix A Annotation Guide
A.1 Labels
@@
The name of a company or financial institution mentioned in the agreement. Examples: "Goldman Sachs", "JP Morgan Chase".
@@
The role that an organization holds in the agreement. Examples: "Lender", "Borrower", "Guarantor".
@@
A more specific classification within an Organization Role that provides additional detail. Examples: "Lead Lender", "Administrative Agent".
@@
The name of an individual explicitly mentioned in the agreement. Examples: "John Doe", "Jane Smith".
@@
The title or role held by a person within an organization. Examples: "Vice President", "Loan Officer".
@@
A physical address associated with an organization. Examples: "35 Peachtree St, Atlanta, GA".
@@
A classification of a Location, providing additional information on its function. Example: "Headquarters", "Branch Office".
A.2 General rules
@@
Example: In Figure 5(a), "BANK OF AMERICA, N.A." appears multiple times but is annotated only once. However, its roles, "Administrative Agent" and "Lender", are each annotated where they are explicitly stated. If another organization is a lender and/or an administrative agent, those role labels will be annotated as close as possible to the name label. Similarly for "Alexandra M. Knights" and her position "Authorized Signatory".
@@
  • Org Name $\rightarrow$ Org Role: Each organization name is annotated once per document, while its role is annotated as close to the name as possible. If multiple organizations share the same role (e.g., Borrower) and it is mentioned collectively in the signature box, the role is annotated once and linked to all relevant organizations. Otherwise, each occurrence of a role is annotated separately and linked to the corresponding organization.
  • Org Role $\rightarrow$ Org Sub-Role: If an organization has a specific sub-role (e.g., Parent Borrower), its role is annotated separately from others, and the sub-role is linked to it. The connection follows a hierarchical structure, linking the organization to its role, then the role to its sub-role.
  • Org Name $\rightarrow$ Person Name: Each person’s name is annotated once per document and linked to a single organization (see A.3 for special scenarios).
  • Person Name $\rightarrow$ Person Position: A person’s position is annotated as close to their name as possible. Individuals may hold multiple positions, which are all linked to the same person annotation. In some cases, position mentions may be missing (see Figure 5(d) and A.3).
  • Org Name $\rightarrow$ Location: Locations are annotated once per document and linked to the relevant organization.
  • Location $\rightarrow$ Location Type: Each location has exactly one type, which is annotated as close to the location name as possible. Location types are often missing, and in such cases, no annotation is applied (see Fig 5(d)).
@@
In cases where the text to be annotated is split across multiple lines or interrupted by symbols, preventing a single coherent annotation, we annotate each segment separately and link them sequentially: the first to the second, the second to the third, and so on. This issue is particularly common for Locations, where addresses or names are fragmented (see Figure 5(d)). Since our annotation structure does not include relations between labels of the same type, these links are easily identifiable. During post-processing, entities connected by continuation links are automatically merged with spaces, ensuring a single coherent label.
A.3 Main Edge Cases
@@
One of the extracted credit agreements has only 2 organizations involved, and their roles are mentioned neither in the signature box nor in the header (Figure 5(f)). Moreover, the role of the lender is not mentioned anywhere in the document, while the role of the borrower is mentioned once (Figure 5(g)).

### v2

@@ 3.1 Baselines
We evaluate 16 proprietary and open-weight LLMs on KG-MuLQA-D benchmark.
@@
Gemini-2.0-Flash & Gemini-2.0-Flash-Lite (Team et al., 2024a), Gemini-1.5-Pro & Gemini-1.5-Flash (Team et al., 2023, 2024a), and GPT-4o & GPT-4o-Mini (Achiam et al., 2023; Hurst et al., 2024).
Open-Weight LLMs:
DeepSeek-R1 (DeepSeek-AI et al., 2025), DeepSeek-V3 (Liu et al., 2024a), GPT OSS 120B (OpenAI et al., 2025), Llama-4-Maverick-17B-128E-Instruct-FP8 (AI, 2024a), Llama-4-Scout-17B-16E-Instruct (AI, 2024b), Llama-3.1-405B-Instruct-Turbo & Llama-3.1-70B-Instruct & Llama-3.1-8B-Instruct (Dubey et al., 2024), Qwen3-235B-Instruct (Yang et al., 2025), Qwen2-72B-Instruct (Yang et al., 2024).
Given that our documents average 80k tokens, we restrict our evaluation to models supporting a context length of at least 128k tokens. Nonetheless, even long-context models require a carefully designed prompting strategy to extract relevant answers effectively (see Section D.1). We made several attempts to evaluate Gemma-2-8k (Team et al., 2024b), Mixtral-Small-32k, and Mixtral-8×22-65k (Jiang et al., 2024a), but these models returned “Not found” in the vast majority of the cases, making systematic evaluation infeasible.
For proprietary models and those accessed via Together.AI [^5] https://www.together.ai/, we utilized the LangChain [^6] https://www.langchain.com/ framework, while open-weight models were employed using vLLM (Kwon et al., 2023) for efficient inference. We set the temperature to 0.0 to support reproducibility of our results.
@@ 3.2 Experiment Setup
Figure 3: This figure illustrates the multi-stage process used to evaluate long-context LLMs: documents are split into chunks (if exceeding 128K tokens), paired with question batches ($\leqslant$50), and sequentially fed to the model. Model responses are collected per chunk and per question batch to ensure evaluation scalability across large documents.
Our evaluation setup is designed to handle long financial documents and large batches of questions in a way that fits the context length constraints of modern language models. As seen in Figure 3, we split a document into chunks if it exceeds 128K tokens: most of the evaluated models have context length of 128K tokens, so it is impossible to evaluate these models without chunking. For the larger models, we have to keep the same chunking strategy to ensure a fair comparison between models (see Appendix D.5 for the effect of chunking on evaluation). Then, we prompt the model with the document (or chunk) and a batch of 50 questions associated with it: long-context evaluation is very costly, so batching allows at most 3 inference runs per document (given almost 150 questions for some documents). This design ensures consistent and scalable evaluation across documents of varying lengths (see Appendix D.1 for further details).
We use four evaluation metrics: F1 score, Edit Distance, Cosine Similarity, and LLM-as-a-Judge. Among these, F1 and LLM-as-a-Judge are highlighted in Table 4 for their complementary strengths: F1 quantifies exact token-level overlap, while LLM-J captures semantic correctness with human-aligned scoring (see Appendix D.2 for details). Both metrics show the highest correlation with human evaluation metrics (see Appendix D.3).
@@ 3.3 Main Results
As seen in Table 4, the results for simple questions are straightforward to interpret: more advanced models mostly get better scores. For medium questions, DeepSeek-V3 outperforming DeepSeek-R1 and Gemini-1.5-Pro outperforming Gemini-2.0 mainly comes from fewer abstentions and “Not Found” cases, which raises recall and F1 but doesn’t reflect stronger reasoning. Llama-3.1 405B behaves similarly: it answers more often, boosting both metrics through coverage rather than logic. In contrast, Qwen2’s high judge score but low F1 reflects a different pattern: it tends to produce verbose, loosely correct text that seems semantically reasonable, so the judge rates it highly even when exact matches are wrong or incomplete. The GPT OSS 120B, however, gets good performance on medium difficulty from its ability to extract entities while minimizing errors, as it strikes a balance between answering and abstaining.
For hard questions on proprietary models, Gemini-2.0-Flash shows high judged quality but lower F1, while Flash-Lite reverses this trend; Gemini-1.5-Pro remains balanced. This difference comes from how often each model answers: more attempts raise F1 through partial overlaps with the gold, while fewer but cleaner attempts earn better judged quality (given responses are not verbose).
For open-weight models, Llama-4-Maverick achieves a high LLM-J score with moderate F1 for answering selectively with short, factual fragments that look semantically right but miss full token matches. Llama-3.1 405B, by contrast, produces longer and more complete answers that align better with references, yielding higher F1 while maintaining strong judged coherence, hence demonstrating the best performance with hard questions.
Overall, as task complexity increases, models split between answering broadly and answering cautiously. This divergence weakens the correlation between F1 and judge scores, since literal overlap and perceived semantic adequacy start reflecting different behaviors.
4 Analysis
Table 4: This table presents the performance of 16 LLMs, evaluated across Easy, Medium, and Hard question categories. The metrics include the F1 Score and the LLM-as-a-Judge rating, capturing both token-level accuracy and semantic correctness (see the appendix D.2 for extended evaluation). The results reveal a consistent decline in performance as question complexity increases, with notable model-specific strengths and weaknesses. denotes the models evaluated on a smaller subset due to cost constraints (see Appendix D.4 for details).
@@ 4 Analysis
Open-weight LLMs
DeepSeek-R1 685B 128K 0.6269 3.6965 0.3293 2.6354 0.2609 2.0736
@@ 4.1 Overall Error Trends
As question complexity increases, the LLM’s ability to retrieve and generate correct responses degrades markedly. With increase in number of hops or set operations, the percentage distribution of each error type increases (see Figure 4 for details).
@@ 4.2 Error Analysis
Misinterpretation of Semantics LLMs frequently misinterpret task-specific terms such as “location type” or “position”. Rather than grounding responses in document context, models often default to general world knowledge (e.g., answering “city” for a location type), indicating weak comprehension of domain-specific semantics.
Implicit Information Gaps The model struggles to extract answers that are not explicitly stated but implied through document structure, such as signature blocks or formatting cues. For instance, it fails to associate a person with an organization unless the relationship is stated verbatim, ignoring visual or positional indicators that a human reader would readily infer. We consider it a potential reason for why models perform slightly better on full documents than in the oracle setting for hard questions (see Appendix D.8 for details).
Set Operation Failures To examine the impact of the number of set operations on model performance, we conducted an ablation experiment (Figure 11(a)) by varying one dimension within our three-dimensional parameter space while holding the others constant. Specifically, we increased the number of set operations while keeping the plurality and hop dimensions fixed. We observe that an increase in the number of set operations consistently resulted in a decline in performance metrics. Questions involving comparisons, intersections, or exclusions across entities consistently yield low F1 scores. The model fails to isolate shared or distinct roles when multiple organizations or individuals are involved, indicating limited capacity for multi-hop or set-based reasoning.
Long-Context Retrieval Errors Despite the presence of answers in the document, models often return “Not Found”, especially when relevant information appears in the middle of the context. This aligns with the “lost-in-the-middle” phenomenon (Liu et al., 2024b), where models disproportionately focus on the beginning or end of long texts.
@@
5 Related Work
[…]
Most existing benchmarks do not define explicit question complexity levels, treating all QA instances as uniformly difficult. Several recent datasets (Gupta et al., 2024; Zhao et al., 2024c; Zhu et al., 2024; Hsieh et al., 2024; Kuratov et al., 2024) vary task difficulty by changing task type or length, distributing supporting facts in long contexts, or altering context length and needle placement. We introduce a granular notion of question complexity, defined along multiple dimensions (multi-hop reasoning, set operations, and answer plurality), allowing systematic control over difficulty and enabling fine-grained error analysis (Section 4.2) that reveals distinct model failure modes. While MuSiQue also organizes questions by reasoning depth (2–4 hops), it lacks the multi-dimensional decomposition we propose. Our formulation extends beyond hop count to capture qualitatively different reasoning types, which provides a richer understanding of model strengths and weaknesses.
Figure 4: Trends in error types across complexity levels. The template level is given by P+H+#SOP+H+\#SO.
@@ 6 Discussion & Conclusion
Our work introduces KG-MuLQA, a framework that programmatically constructs long-context QA benchmarks using knowledge-graph representations of real-world documents. We begin by annotating credit agreements with a structured schema to capture entities and relations, convert them into RDF-based knowledge graphs, and generate multi-level QA pairs via deterministic SPARQL templates that vary in hop count, set operations, and answer plurality. We then evaluate 16 long-context LLMs across these controlled complexity levels using 4 complementary metrics and perform human evaluation to validate metric alignment. Our quantitative and qualitative analyses reveal consistent weaknesses in multi-hop retrieval, set-based reasoning, and implicit relation grounding, establishing KG-MuLQA as a scalable and interpretable framework for assessing long-context comprehension.
Human Performance Evaluation
Appendix D.7 shows human performance remains stable across difficulty. This baseline both grounds our metric choices and indicates model errors stem from reasoning/retrieval limits rather than dataset artifacts.
@@
As discussed in Appendix A.1, the ability to process long documents is essential for real-world QA tasks. Our benchmark illustrates this challenge using credit agreements, where information is densely structured, scattered, and cross-referenced. Legal contracts, policy documents, patent filings, and other non-legal such as medical records or scientific articles all exhibit similarly complex entity structures, making our methodology broadly applicable. To illustrate the applicability and generalizability of our framework, we performed a full-scale ablation study (see Appendix F) on medical documents, including defining the annotation schema, constructing question templates, etc.
Limitations
The created QA pairs are based on financial credit agreements only. As with most existing QA benchmarks, the benchmark favors deductive logic because the dataset includes graph-representable questions.
Ethical Considerations
We generally avoid mentioning specific agreements and their details. Instead, we use placeholders such as ABCDE Inc. for organization names, Richard Roe for person names, and 10 Peachtree... for locations to illustrate the structure of agreement entities and relations. However, in our annotation guide (Appendix A) and error analysis (Appendix E), we include screenshots of real document annotations and mention specific examples to highlight the complexity of these agreements. Moreover, we store the raw HTML files of the agreements in our GitHub repository, enabling other researchers to reproduce our benchmark or conduct further experiments. As these documents are publicly available through the SEC’s EDGAR system, their inclusion for both illustrative and experimental purposes raises no ethical concerns.
@@ Appendix A Annotation Guide
A.1 Suitability of Credit Agreements
Credit agreements are particularly well-suited for long-context evaluation because they are rich in entities and relations, involving numerous organizations, individuals, roles, and locations that appear in varied forms throughout the document, making both retrieval and reasoning challenging. They also exhibit ambiguous co-reference, as entities may be referred to inconsistently (e.g., “Bank of America, N.A.” vs. “Bank of America”), demanding precise resolution of abbreviations and partial mentions. While structured but not standardized, these documents follow broad templates (e.g., loan terms, covenants, defaults) but vary greatly in phrasing and layout, preventing reliance on fixed patterns and requiring genuine comprehension. Finally, their real-world significance as legally binding records of multi-million-dollar transactions makes accurate understanding essential for assessing risk, ensuring compliance, and supporting financial and regulatory decisions. Supporting examples illustrating these aspects are provided in the latter subsections of this Appendix.
A.2 Labels
@@
The name of a company or financial institution mentioned in the agreement. Examples: “Goldman Sachs”, “JP Morgan Chase”.
@@
The role that an organization holds in the agreement. Examples: “Lender”, “Borrower”, “Guarantor”.
@@
A more specific classification within an Organization Role that provides additional detail. Examples: “Lead Lender”, “Administrative Agent”.
@@
The name of an individual explicitly mentioned in the agreement. Examples: “John Doe”, “Jane Smith”.
@@
The title or role held by a person within an organization. Examples: “Vice President”, “Loan Officer”.
@@
A physical address associated with an organization. Examples: “35 Peachtree St, Atlanta, GA”.
@@
A classification of a Location, providing additional information on its function. Example: “Headquarters”, “Branch Office”.
A.3 General rules
@@
Example: In Figure 5(a), “BANK OF AMERICA, N.A.” appears multiple times but is annotated only once. However, its roles, “Administrative Agent” and “Lender”, are each annotated where they are explicitly stated. If another organization is a lender and/or an administrative agent, those role labels will be annotated as close as possible to the name label. Similarly for “Alexandra M. Knights” and her position “Authorized Signatory”.
@@
[…]
  • $\rightarrow Location$ Type: Each location has exactly one type, which is annotated as close to the location name as possible. Location types are often missing, and in such cases, no annotation is applied (see Fig 5(d)).
@@
In cases where the text to be annotated is split across multiple lines or interrupted by symbols, preventing a single coherent annotation, we annotate each segment separately and link them sequentially: the first to the second, the second to the third, and so on. This issue is particularly common for Locations, where addresses or names are fragmented (see Figure 5(d)). Since our annotation structure does not include relations between labels of the same type, these links are easily identifiable. During post-processing, entities connected by continuation links are automatically merged with spaces into a single coherent label.
A.4 Main Edge Cases
@@
One of the extracted credit agreements involves only 2 organizations, and their roles are mentioned neither in the signature box nor in the header (Figure 5(f)). Moreover, the role of the lender is not mentioned in the document, while the role of the borrower is mentioned once (Figure 5(g)).
A.5 Annotation Process Details
We annotated the extracted credit agreements using Label-Studio [^7] https://github.com/HumanSignal/label-studio, supporting entity linking and relation extraction for .html files. The annotation was carried out by annotators with backgrounds in Computer Science or Mathematics, under the direct supervision of a Master of Finance and a PhD student advised by a Chair Professor of Finance, also a co-author. The annotation followed a structured, iterative approach to ensure consistency and resolve edge cases. The team first established label definitions and completed an initial annotation round. The results were reviewed to identify and discuss edge cases (Appendix A.4), which led to refinements in the annotation guidelines and the second round of annotations (inter-annotator agreement was not calculated at this point, as the guidelines were evolving). Following the second annotation round, we constructed knowledge-graph (KG) document representations, sampled graphs, and compared them against the ground-truth annotations. Insights from this qualitative assessment (Appendix B.1) informed a third annotation round, after which we applied an additional cleaning script to the KG representations.
A.6 Complexity vs coverage
KG-MuLQA defines seven entity and six relation types, and this choice reflects a deliberate trade-off between semantic coverage and compositional complexity. Expanding the ontology would increase topical breadth but would not necessarily yield deeper reasoning requirements. Our goal is to evaluate long-context comprehension under settings where information is distributed across distant parts of a document and must be retrieved, composed, and reasoned over.
Thus, the benchmark emphasizes complex inference chains (multi-hop reasoning, set operations, and answer plurality) over the schema coverage. Answering questions from our dataset requires linking dispersed mentions, disambiguating roles, and integrating partial evidence: operations that jointly test both retrieval and reasoning. This structure allows us to probe failure modes of long-context models in a controlled environment, balancing interpretability, scalability, and difficulty without relying on full document coverage.